# A Hierarchical Approach to Quantum Many-Body Systems in Structured Environments

### v1

@@ II.2 Combining HEOM with BBGKY
Our goal in the following is to derive a closed hierarchy of equations for the reduced density matrix $F_{12}$, such that the (auxiliary) density matrices in Eq. (9) only scale with the square of the dimension of the single-particle Hilbertspace. From Eq. (9) it follows that (see App. B)
$\begin{split}\dot{F}_{12}^{(\boldsymbol{n},\boldsymbol{m})}&=-i[H_{1}+H_{2},F_{12}^{(\boldsymbol{n},\boldsymbol{m})}]+\operatorname{Tr}_{3}\left([V_{13}+V_{23},F_{123}^{(\boldsymbol{n},\boldsymbol{m})}]\right)-\left(\boldsymbol{w}\cdot\boldsymbol{n}+\boldsymbol{w}^{*}\cdot\boldsymbol{m}\right)F_{12}^{(\boldsymbol{n},\boldsymbol{m})}\\ &+\sum_{k=1}^{M}\bigg{(}G_{k}\left(n_{k}(L_{1}+L_{2})F_{12}^{(\boldsymbol{n}-\boldsymbol{e}_{k},\boldsymbol{m})}+n_{k}\operatorname{Tr}_{3}(L_{3}F_{123}^{(\boldsymbol{n},\boldsymbol{m})})+m_{k}F_{12}^{(\boldsymbol{n},\boldsymbol{m}-\boldsymbol{e}_{k})}(L_{1}^{\dagger}+L_{2}^{\dagger})+m_{k}\operatorname{Tr}_{3}(L_{3}^{\dagger}F_{123}^{(\boldsymbol{n},\boldsymbol{m})})\right)\\ &\quad\quad+\left[F_{12}^{(\boldsymbol{n}+\boldsymbol{e}_{k},\boldsymbol{m})},L_{1}^{\dagger}+L_{2}^{\dagger}\right]+\left[L_{1}+L_{2},F_{12}^{(\boldsymbol{n},\boldsymbol{m}+\boldsymbol{e}_{k})}\right]\bigg{)}.\end{split}$ (11)
@@
For any interacting many-body system, the two-body RDM $F_{12}$ will couple to the three-body RDM $F_{123}$, which in turn will couple to the four-body RDM and so on. In order to close this hierarchy we therefore need to express the three-body RDM in terms of the one- and two-body RDM
$F_{123}\approx\widetilde{F}_{123}(F_{12},F_{1}).$
@@
We demonstrate in App. B how this can be achieved by reformulating the hierarchy of equations for the density matrices $\rho_{sys}^{(\boldsymbol{n},\boldsymbol{m})}$ as an equation for a single operator in an extended Hilbertspace $\mathcal{H}_{ext}=\mathcal{H}_{sys}\otimes\mathcal{H}_{aux}$ and subsequently neglecting three-body correlations within the system. However, we keep correlations between two particles in the system and the auxiliary degree of freedom that captures the influence of the environment. With these approximations we arrive at the expression below for the three-body RDM, where we have left out indices $(\boldsymbol{0},\boldsymbol{0})$ for brevity
$\begin{split}\widetilde{F}^{(\boldsymbol{m},\boldsymbol{n})}_{123}&=4\frac{(N-1)(N-2)}{N^{3}}\operatorname{Tr}(F_{1}^{(\boldsymbol{m},\boldsymbol{n})})F_{1}F_{2}F_{3}\\ &+\frac{N-2}{N}\left(F_{12}F_{3}^{(\boldsymbol{m},\boldsymbol{n})}+F_{2}^{(\boldsymbol{m},\boldsymbol{n})}F_{13}+F_{1}^{(\boldsymbol{m},\boldsymbol{n})}F_{23}\right)\\ &+\frac{N-2}{N}\left(F_{1}F_{23}^{(\boldsymbol{m},\boldsymbol{n})}+F_{13}^{(\boldsymbol{m},\boldsymbol{n})}F_{2}+F_{12}^{(\boldsymbol{m},\boldsymbol{n})}F_{3}\right)\\ &-\frac{N-2}{N^{2}}\operatorname{Tr}(F_{1}^{(\boldsymbol{m},\boldsymbol{n})})\left(F_{12}F_{3}+F_{13}F_{2}+F_{1}F_{23}\right)\\ &-2\frac{(N-2)(N-1)}{N^{2}}\Big{(}F_{1}^{(\boldsymbol{m},\boldsymbol{n})}F_{2}F_{3}\\ &\quad+F_{1}F_{2}^{(\boldsymbol{m},\boldsymbol{n})}F_{3}+F_{1}+F_{2}F_{3}^{(\boldsymbol{m},\boldsymbol{n})}\Big{)}.\end{split}$ (12)
@@
Inserting the above approximation into Eq. (11) provides the BBGKY-HEOM equation up to three-body correlations – the central tool of this work. As desired, it represents a closed hierarchy for the two-particle RDM that can account for interactions within the system and simultaneously includes the influence of a generic non-Markovian environment.
@@
$\widetilde{F}^{(\boldsymbol{m},\boldsymbol{n})}_{123}$ is invariant under particle exchange, but does not preserve the (anti-)symmetry of the wavefunctions as required for (fermionic) bosonic particles. As explained in detail in App. B this can be achieved by employing an (anti-) symmetrization operator which projects $\widetilde{F}^{(\boldsymbol{m},\boldsymbol{n})}_{123}$ onto the (anti-) symmetric subspace of the the Hilbertspace while retaining its norm [50].
@@ III Results
Every quantum system is a collection of individual particles, may it be a collection of single-photon emitters or the individual electronic and vibrational degrees of freedom that constitute one such emitter. We demonstrate in the following the applicability, strength, and limitation of BBGKY-HEOM for both of these cases. In particular, we focus on two well known paradigmatic models from quantum optics and condensed matter physics: the driven Tavis-Cummings and Fermi-Hubbard model.
Figure 1: Squeezing and superradiance in the driven Tavis-Cummings model: A) Spin dynamic for $N=50$, $g\sqrt{N}/\kappa=0.5$, $\Delta_{z}/\kappa=0.2$, $\Omega/\kappa=0.05$, $\Delta_{c}/\kappa=10$. B) Corresponding spin-squeezing (see text). C) Superradiant emission from fully excited state. Parameters adjusted to $\Delta_{z}/\kappa=2$, $\Delta_{x}/\kappa=1$, $\Delta_{c}/\kappa=2.4$, $g/\kappa=0.1$. D) Photon-occupation normalized by $N^{2}$ over time and maximum photon number in burst (not normalized) for different N obtained with BBGKY-HEOM. BBGKY-HEOM provides reliable predictions and only begins to deviate slightly at long times when explicit 3-body correlations accumulate to noticeable values.
@@ III.1 Driven-dissipative many-emitter systems
A collection of emitters that share a common mode can interact with it collectively, potentially resulting in burst-like correlated emission known as superradiance. Such phenomena are closely linked to the non-classical dynamic of squeezed spin states. Following a series of simplifications for matter, cavity, and light-matter coupling, one obtains the widely used Dicke Hamiltonian [53]
@@
The Pauli matrices $\sigma_{x,y,z}^{i}$ represent the SU(2) algebra of an idealized two-level emitter while $a^{\dagger}$($a$) are the creation (annihilation) operators of the cavity mode. The collection of emitters can be brought out of equilibrium by continuously pumping energy into the system $H_{drive}=2\Omega\cos{\omega_{d}t}\sum_{i}\sigma_{x}^{i}$. A unitary transformation moves us into a frame that rotates with the driving frequency $\omega_{d}$. As long as the drive is close to resonance, i.e., $\omega_{d}\gg\Delta_{z},\Delta_{x},g,\Omega$ with the detunings $\Delta_{z}=\omega_{a}-\omega_{d}$, $\Delta_{c}=\omega_{c}-\omega_{d}$, we can safely discard counter-rotating terms $e^{i(\omega_{d}+\omega_{a/c})}$ that oscillate too fast to influence the dynamic of the system. We obtain the paradigmatic driven Tavis-Cummings Hamiltonian
@@
We will consider the emitters or spins in the following as our system of interest and the cavity as environment dissipating energy into free space at a rate $\kappa$. The bath correlation function is readily obtained, for example from the Heisenberg-Langevin equations [51], as
@@
Given the necessary computational resources, we could have just as well decided to directly solve the entire system of N-emitter plus lossy cavity using the Lindblad master equation
@@
This exact solution will be used to estimate the quality of the BBGKY-HEOM and can be obtained in a collective spin picture $S_{k}=\sum_{i}\sigma_{k}^{i}$, $k\in\{x,y,z,+,-\}$. The $2^{N}$ dimensional system Hilbert space can then be restricted to a basis of $N+1$ permutationally symmetric Dicke states {|$\{\ket{N/2,m},m=-N/2,-N/2+1,\dots,N/2\}$ which results in a linear increase of the relevant basis size with N and therefore a quadratic increase of the Liouville space. However, the collective spin-picture fails as soon as individual emitter-emitter interactions need to be included, resulting in an exponentially growing Hilbert space that prohibits exact solutions for systems larger than a handful of emitter. The computational cost for BBGKY-HEOM on the other hand is unaffected by this change and remains system-size independent. Note, that even if there is no direct interaction in this model the spins still interact indirectly via the cavity mode.
Figure 2: Fermi-Hubbard model coupled to lossy cavity mode: (Left) Dynamic of dipole moment of Fermi-Hubbard model ($J=1$, parameters in text) after site-dependent quench V(t) = $\sum_{i=0}^{3}V_{i}(t)c_{i}^{\dagger}{c}_{i}$ with $V_{i}(0^{+})=\frac{1}{5(i+1)}$, which mimics a charge localization event, for different onsite repulsion $U$. BBGKY-HEOM (blue solid) provides reliable predictions for weak to moderate onsite repulsion when compared to the exact solution (black dashed). (Right) Dipole moment driven by time-dependent potential $V_{i}(t)=\frac{f(t)}{5(i+1)}$, where f(t) is a step function as shown by the grey dashed line and $U=0.1$. The example simulates charge noise due to the movement of electrons in the vicinity of the molecule. Inset illustrates onsite occupation.
@@
Fig. 1 (A) illustrates the system dynamic in the rotating frame for $N=50$ emitter strongly coupled to the cavity $\sqrt{N}g/\kappa=\sqrt{N/4}$ in comparison to the exact solution of the Master equation. The spin components $\langle S_{i}\rangle$ are accurately predicted within the simulation window, despite the non-classical dynamic of the system. Shown in the upper inset is the spin squeezing (see App. A) which is exhibiting values smaller than unity – a sufficient condition for entanglement. We observes small deviations at long times which can be understood by investigating the slow increase in the explicit 3-body correlations (lower inset), i.e., the term that is discarded in the currently chosen truncation. The spin-squeezing is further illustrated by the spin-Q function $Q=\bra{\theta,\phi}\rho\ket{\theta,\phi}$ in the upper right subplot (B), with |$\ket{\theta,\phi}$ the spin coherent state. Mean-field theory will fail in predicting any of the spin-squeezing effects.
@@
Let us fully excite all emitters and illustrate the resulting superradiant emission in Fig. 1 (C) (parameters adjusted, see caption). BBGKY-HEOM follows closely the exact dynamic and small deviations are visible only close to the fully de-excited steady-state. While mean-field theory models the initial frequency quite accurately, it fails to model the spontaneous emission process and remains trapped at the upper fix-point. The inset projects this dynamic on the Bloch-sphere which illustrates that the failure of the mean field can be connected to restriction to the surface of the sphere while the correct dynamic requires to transverse it diagonally. The sudden increase in the photon number originating from the burst-like emission of $N$ correlated emitter is presented in Fig. 1 (D), demonstrating again accurate predictions by BBGKY-HEOM. Continuous driving results in a continuous replenishment of the cavity mode. The maximum photon-number (inset) is correctly described in its quadratic increase with the emitter number $N$. In conclusion, BBGKY-HEOM provides accurate predictions for matter and cavity observables at constant cost, i.e., for arbitrary many emitter and with the option to account for direct interaction. We will now relax the simplification of the emitter and instead combine the dynamic of the many-body correlated system with a structured system-bath coupling.

### v2

@@ II.2 Combining HEOM with BBGKY
Our goal in the following is to derive a closed hierarchy of equations for the reduced density matrix $F_{12}$, such that the (auxiliary) density matrices in Eq. (II.2) only scale with the square of the dimension of the single-particle Hilbert space. From Eq. (II.2) it follows that (see App. C)
$\dot{F}_{12}^{(\bm{n},\bm{m})}$ $=-i[H_{1}+H_{2},F_{12}^{(\bm{n},\bm{m})}]-i\operatorname{Tr}_{3}\left([V_{13}+V_{23},F_{123}^{(\bm{n},\bm{m})}]\right)-\left(\bm{w}\cdot\bm{n}+\bm{w}^{*}\cdot\bm{m}\right)F_{12}^{(\bm{n},\bm{m})}$
$+\sum_{k=1}^{M}\bigg{(}G_{k}\left(n_{k}(L_{1}+L_{2})F_{12}^{(\bm{n}-\bm{e}_{k},\bm{m})}+n_{k}\operatorname{Tr}_{3}(L_{3}F_{123}^{(\bm{n}-\bm{e}_{k},\bm{m})})\right)+G_{k}^{*}\left(m_{k}F_{12}^{(\bm{n},\bm{m}-\bm{e}_{k})}(L_{1}^{\dagger}+L_{2}^{\dagger})+m_{k}\operatorname{Tr}_{3}(L_{3}^{\dagger}F_{123}^{(\bm{n},\bm{m}-\bm{e}_{k})})\right)$
$\quad\quad+\left[F_{12}^{(\bm{n}+\bm{e}_{k},\bm{m})},L_{1}^{\dagger}+L_{2}^{\dagger}\right]+\left[L_{1}+L_{2},F_{12}^{(\bm{n},\bm{m}+\bm{e}_{k})}\right]\bigg{)}.$ (12)
@@
For any interacting many-body system, the two-body RDM $F_{12}$ will couple to the three-body RDM $F_{123}$, which in turn will couple to the four-body RDM and so on. In order to close this hierarchy we express the three-body RDM in terms of the one- and two-body RDM
$F_{123}\approx\widetilde{F}_{123}(F_{12},F_{1}),$
@@
where $\widetilde{F}_{123}$ should be understood as a function that approximates the true three-body RDM. We demonstrate in App. C how this can be achieved by reformulating the hierarchy of equations for the density matrices $\rho_{sys}^{(\bm{n},\bm{m})}$ as an equation for a single operator in an extended Hilbert space $\mathcal{H}_{ext}=\mathcal{H}_{sys}\otimes\mathcal{H}_{aux}$ and subsequently neglecting three-body correlations within the system. However, we keep correlations between two particles in the system and the auxiliary degree of freedom that captures the influence of the environment. With these approximations we arrive at the expression below for the three-body RDM, where we have left out indices $(\bm{0},\bm{0})$ for brevity
$\widetilde{F}^{(\bm{n},\bm{m})}_{123}$ $=4\frac{(N-1)(N-2)}{N^{3}}\operatorname{Tr}(F_{1}^{(\bm{n},\bm{m})})F_{1}F_{2}F_{3}$
$+\frac{N-2}{N}\left(F_{12}F_{3}^{(\bm{n},\bm{m})}+F_{2}^{(\bm{n},\bm{m})}F_{13}+F_{1}^{(\bm{n},\bm{m})}F_{23}\right)$
$+\frac{N-2}{N}\left(F_{1}F_{23}^{(\bm{n},\bm{m})}+F_{13}^{(\bm{n},\bm{m})}F_{2}+F_{12}^{(\bm{n},\bm{m})}F_{3}\right)$
$-\frac{N-2}{N^{2}}\operatorname{Tr}(F_{1}^{(\bm{n},\bm{m})})\left(F_{12}F_{3}+F_{13}F_{2}+F_{1}F_{23}\right)$
$-2\frac{(N-2)(N-1)}{N^{2}}\Big{(}F_{1}^{(\bm{n},\bm{m})}F_{2}F_{3}$
$\quad+F_{1}F_{2}^{(\bm{n},\bm{m})}F_{3}+F_{1}+F_{2}F_{3}^{(\bm{n},\bm{m})}\Big{)}.$ (13)
@@
Inserting the above approximation into Eq. (II.2) provides the BBGKY-HEOM equation up to three-body correlations – the central tool of this work. As desired, it represents a closed hierarchy for the two-particle RDM that can account for interactions within the system and simultaneously includes the influence of a generic non-Markovian environment.
@@
$\widetilde{F}^{(\bm{m},\bm{n})}_{123}$ is invariant under particle exchange, but does not preserve the (anti-)symmetry of the wavefunctions as required for (fermionic) bosonic particles. As explained in detail in App. C this can be achieved by employing an (anti-) symmetrization operator which projects $\widetilde{F}^{(\bm{m},\bm{n})}_{123}$ onto the (anti-) symmetric subspace of the Hilbert space while retaining its norm [66].
@@ III Results
Every quantum system is a collection of individual particles, may it be a collection of single-photon emitters or the individual electronic and vibrational degrees of freedom that constitute one such emitter. In the following, we demonstrate the applicability, strength, and limitation of BBGKY-HEOM for both of these cases. In particular, we focus on three systems from quantum optics and condensed matter physics. Benchmarks via driven Tavis-Cummings and Cavity-Fermi-Hubbard models serve to demonstrate the excellent performance of BBGKY-HEOM. We then go beyond the state of the art with applications to few-emitter molecular lasing and system-bath competition in the Cavity-Fermi-Hubbard model.
Figure 1: Squeezing and superradiance in the driven Tavis-Cummings model: (a) Spin dynamics for $N=50$, $g\sqrt{N}/\kappa=0.5$, $\Delta_{z}/\kappa=0.2$, $\Omega/\kappa=0.05$, $\Delta_{c}/\kappa=10$. (b) Corresponding spin-squeezing (see text). (c) Superradiant emission from fully excited state. Parameters adjusted to $\Delta_{z}/\kappa=2$, $\Delta_{x}/\kappa=1$, $\Delta_{c}/\kappa=2.4$, $g/\kappa=0.1$. (d) Photon-occupation normalized by $N^{2}$ over time and maximum photon number in burst (not normalized) for different N obtained with BBGKY-HEOM. BBGKY-HEOM provides reliable predictions and only begins to deviate slightly at long times when explicit 3-body correlations accumulate to noticeable values.
@@ III.1 Driven-dissipative many-emitter systems
A collection of emitters that share a common mode can interact with it collectively, potentially resulting in burst-like correlated emission known as superradiance and non-classical squeezed spin states. Following a series of simplifications for matter, cavity, and light-matter coupling, one obtains the widely used Dicke Hamiltonian [71]
@@
The Pauli matrices $\sigma_{x,y,z}^{i}$ represent the SU(2) algebra of an idealized two-level emitter while $a^{\dagger}$($a$) are the creation (annihilation) operators of the cavity mode. The collection of emitters can be brought out of equilibrium by coherently driving the system $H_{drive}=2\Omega\cos{\omega_{d}t}\sum_{i}\sigma_{x}^{i}$. A unitary transformation moves us into a frame that rotates with the driving frequency $\omega_{d}$. As long as the drive is close to resonance, i.e., $\omega_{d}\gg\Delta_{z},\Delta_{x},g,\Omega$ with the detunings $\Delta_{z}=\omega_{a}-\omega_{d}$, $\Delta_{c}=\omega_{c}-\omega_{d}$, we can safely discard counter-rotating terms $e^{i(\omega_{d}+\omega_{a/c})}$ that oscillate too fast to influence the dynamics of the system. We obtain the paradigmatic driven Tavis-Cummings Hamiltonian
@@
We will consider the emitters or spins in the following as our system of interest and the cavity as an environment that dissipates energy into free space at a rate $\kappa$. The bath correlation function is readily obtained, for example, from the Heisenberg-Langevin equations [67], as
@@
Given the necessary computational resources, we could have just as well decided to directly solve the entire system of $N$-emitters plus lossy cavity using the Lindblad master equation
@@
This exact solution will be used to estimate the quality of the BBGKY-HEOM and can be obtained in a collective spin picture $S_{k}=\sum_{i}\sigma_{k}^{i}$, $k\in\{x,y,z,+,-\}$. The $2^{N}$ dimensional system Hilbert space can then be restricted to a basis of $N+1$ permutationally symmetric Dicke states {|$\{\ket{N/2,m},m=-N/2,-N/2+1,\dots,N/2\}$ which results in a linear increase of the relevant basis size with $N$ and therefore a quadratic increase of the Liouville space. The collective spin-picture fails as soon as individual emitter-emitter interactions need to be included, resulting in an exponentially growing Hilbert space that prohibits exact solutions for systems larger than a handful of emitters. However, the computational cost for BBGKY-HEOM is unaffected by this change and remains system-size independent. Even if there is no direct interaction in this model, the spins still interact indirectly via the cavity mode.
@@
Fig. 1 (a) illustrates the system dynamics of Eq. (17) for $N=50$ emitters strongly coupled to the cavity $\sqrt{N}g/\kappa=\sqrt{N/4}$ in comparison to the exact solution of the master equation. The spin components $\langle S_{i}\rangle$ are accurately predicted within the simulation window, despite the non-classical dynamics of the system. Shown in the upper inset is the spin squeezing (see App. A) which is exhibiting values smaller than unity – a sufficient condition for entanglement. We observe small deviations at long times which can be understood by investigating the slow increase in the explicit 3-body correlations (lower inset), i.e., the term that is discarded in the currently chosen truncation. The spin-squeezing is further illustrated by the spin-Q function $Q=\bra{\theta,\phi}\rho\ket{\theta,\phi}$ in the upper right subplot (b), with |$\ket{\theta,\phi}$ the spin coherent state. Mean-field theory will fail in predicting any of the spin-squeezing effects.
@@
Let us fully excite all emitters and illustrate the resulting superradiant emission in Fig. 1 (c) (parameters adjusted, see caption). BBGKY-HEOM is accurate, with small deviations to the exact solution only at long times near the fully de-excited steady-state. While mean-field theory models the initial frequency quite accurately, it fails to model the spontaneous emission process and remains trapped at the upper fix-point. The inset projects this dynamic on Bloch’s sphere. The failure of mean field originates from restricting the dynamics to the surface of the sphere, whereas the correct dynamics traverses the Bloch ball diagonally. The sudden increase in the photon number originating from the burst-like emission of $N$ correlated emitter is presented in Fig. 1 (d), demonstrating again accurate predictions by BBGKY-HEOM. Continuous driving results in a continuous replenishment of the cavity mode. The maximum photon-number (inset) is correctly described in its quadratic increase with the emitter number $N$. In conclusion, BBGKY-HEOM provides accurate predictions for matter and cavity observables at constant cost, i.e., for arbitrary many emitter and with the option to account for direct interaction.
III.2 Few-emitter lasing
Figure 2: Overview of the few emitter lasing system.(a) Sketch of the plasmonic nanocavity, comprised of a gold nanoparticle above a gold surface. (b) We restrict our model of the electronic states to the first bright transition, leading to a two level system. However, we take the full spectrum of vibrational modes into account, which we approximate as a large set of harmonic oscillators (purple). Each oscillator in turn couples to a continuum of vibrational modes that we model with an ohmic spectral density (blue). All the vibrations can be combined into a single effective bath with the spectral density $J_{eff}(\omega)$ shown in (d) (green solid line). The corresponding bath correlation function shown in the inset is then fitted with exponentials (black dahed line). Graphic (c) shows a sketch of the lasing mechanism.
For our next application, we turn to a system where no benchmark calculations are available and where BBGKY-HEOM delivers new insights for systems of current experimental relevance. As shown in Fig. 2 (a) a few ($N\approx 2-20$) methylene-blue (MB) molecules are placed inside a plasmonic nanocavity, consisting of an Au nanoparticle on a mirror [79, 80]. Each molecule has a complex internal structure which, in addition to the electronic degrees of freedom, consists of hundreds of internal vibrational modes that in turn couple to a continuum of vibrations in the surrounding dielectric medium. Reference [79] treated this system with the help of simplified effective models, which provided valuable analytic insights and qualitatively captured many experimental features. To achieve these analytical insights, the structure of the vibrational spectrum was simplified to an incoherent drive – a widely applied approximation [81]. However, such simplifications ignore an integral part of molecular physics, limiting understanding and agreement with experiment. We demonstrate in the following how our BBGKY-HEOM method can be used to take the next step and obtain much more detailed results by retaining the vibrational structure. Our approach reveals additional resonant features that are not visible in paradigmatic models.
We restrict the electronic degrees of freedom for each molecule to the many-body states associated with the first bright electronic transition $\widetilde{\omega}_{0}$. The full spectrum of vibrational modes with frequencies $\omega^{vib}_{\lambda}$ and Huang-Rhys factors $S_{\lambda}$ are taken into account. Frequencies and Huang-Rhys factors are estimated using time-dependent density-functional theory (see App. F.1). In the double-harmonic approximation for the electron-phonon coupling, each of the 107 vibrational normal modes is described as a harmonic oscillator, where the equilibrium position of the oscillator is shifted by a displacement $\Delta_{\lambda}=\sqrt{2S_{\lambda}/\omega_{\lambda}}$ if the molecule is in the electronically excited state. Within this approximation, our Hamiltonian for a single molecule takes the form
$H_{mol}=$ $\frac{\widetilde{\omega}_{0}}{2}\sigma^{z}+\sum_{\lambda=1}^{107}\frac{p_{\lambda}^{2}}{2}+\frac{(\omega_{\lambda}^{vib})^{2}}{2}\left(q_{\lambda}-\Delta_{\lambda}\sigma_{+}\sigma_{-}\right)^{2},$
$=$ $\frac{\omega_{0}}{2}\sigma^{z}-\frac{1}{\sqrt{2}}\sum_{\lambda=1}^{107}{\omega_{\lambda}^{vib}}\sqrt{S_{\lambda}}\sigma_{+}\sigma_{-}(b_{\lambda}+b_{\lambda}^{\dagger})$ (18)
$+$ $\sum_{\lambda=1}^{107}\omega_{\lambda}^{vib}b_{\lambda}^{\dagger}b_{\lambda}$
Figure 3: Results for the few emitter lasing system. We compare the qualitative steady state behavior for increasing the coherent drive in the full model (III.2),(20) against an effective incoherent drive that neglects the vibrational structure (III.2). Plots (a) and (b) show the occupation of the excited electronic state, $p_{e}$, and the cavity mode, $a^{\dagger}a$, respectively, for a plasmonic cavity containing ten molecules. Both plots show maxima arising from resonances with the vibrational spectrum that are missed by an effective treatment with an incoherent drive. Those maxima become more pronounced for increasing $N$ as shown in plot (c).
where $q_{\lambda}$, $p_{\lambda}$ are the mass renormalized conjugate position and momenta of the $\lambda$-th normal mode (see Appendix F.1 for a more detailed derivation). The second line reformulated the Hamiltonian in terms of the annihilation (creation) operators $b_{\lambda}$ ($b_{\lambda}^{\dagger}$), with $b_{\lambda}=\sqrt{\omega_{\lambda}/2}(q_{\lambda}+ip_{\lambda}/\omega_{\lambda})$ and we introduced the effective excitation energy $\omega_{0}=\widetilde{\omega}_{0}+\sum_{\lambda}\omega_{\lambda}^{2}\Delta_{\lambda}^{2}/2$, here $\omega_{0}=20857\mathrm{cm}^{-1}$. Due to the embedding of the MB molecules in the surrounding dielectric host medium [80], each of these vibrational modes in turn couples to a continuum of environmental modes, which we describe with an ohmic spectral density (see Fig. 2 (b) and Appendix F.2). These environmental modes lead to a broadening of the vibrational spectrum and ensure that the vibrational modes quickly relax to the corresponding equilibrium position. Due to the rapid relaxation, the setup resembles a four-level lasing scheme in the sense of Fig. 2 (c). Coherent excitation to state |$\ket{4}$ is followed by a fast relaxation into the new equilibrium position of the vibrational modes |$\ket{3}$111 Such a two-step relaxation process of coherent driving combined with vibrational relaxation is often approximated as an effective incoherent drive [81, 79]. This eliminates the presence of vibrational modes, implying a form of wide-band approximation that disregards any structural features. , from which the state decays into |$\ket{2}$ via cavity emission. The decay is then again followed by a fast vibrational relaxation into |$\ket{1}$.
To apply BBGKY-HEOM we proceed by combining all 107 vibrational modes in an exact way into the single (local) bath with effective spectral density [83] $J_{eff}(\omega)=\sum_{\lambda}\lvert g^{eff}_{\lambda}\rvert^{2}\delta(\omega-\omega_{\lambda})$ shown in Fig. 2 (d). This process is described in detail in Appendix F.2.
A few ($N$) of those molecules are now driven by a coherent field with amplitude $2E_{d}$ and frequency $\omega_{0}$. In addition, each molecule is coupled equally with strength $g^{cav}$ to a cavity in resonance with $\omega_{0}$. As detailed in Appendix F.3 we can go into an interaction picture that rotates with the drive frequency $\omega_{0}$ and perform a rotating wave approximation, assuming that $\omega_{0}\gg E_{d},g^{cav}$. The resulting Hamiltonian in the interaction picture then takes the form of Eq. (5):
$H_{tot}^{I}$ $=E_{d}\sum_{j}\sigma_{x}^{j}$
$-\sum_{j}\left(\sum_{\lambda}g_{\lambda}^{eff}\sigma_{+}^{j}\sigma_{-}^{j}(b_{\lambda,j}^{\dagger}+b_{\lambda,j})+\sum_{\lambda}\omega_{\lambda}^{eff}b_{\lambda,j}^{\dagger}b_{\lambda,j}\right)$
$+\sum_{\lambda}g_{\lambda}^{cav}\left(\sum_{j}\sigma_{+}^{j}a_{\lambda}+\sum_{j}\sigma_{-}^{j}a_{\lambda}^{\dagger}\right)$
$+\sum_{\lambda}(\omega_{\lambda}^{cav}-\omega_{0})a_{\lambda}^{\dagger}a_{\lambda}.$ (19)
The second line describes the coupling of each molecule to its local vibrational bath with spectral density $J_{eff}(\omega)$ and the third line describes the coupling to the (global) cavity environment with the Lorentzian spectral density $J_{cav}$. Additionally, we include spontaneous emission according to
$\dot{\rho}_{tot}=-i[H_{tot}^{I},\rho_{tot}]+\Gamma_{\downarrow}\sum_{j}\mathcal{D}[\sigma_{-}^{j}](\rho_{tot}).$ (20)
To apply BBGKY-HEOM we now fit the bath correlation function $\alpha_{eff}(\tau)$ with a sum of 5 exponentials as $\alpha_{eff}(\tau)\approx\sum_{i}G_{i}\exp(-W_{i}\tau)$ as shown in Fig. 2 (d). As the correlation function of the cavity $\alpha_{cav}(\tau)=G_{c}e^{-\kappa\lvert\tau\rvert}$ is already exponential, no additional fit is required. We now propagate the state with different drive strengths $E_{d}<E_{max}=0.1\omega_{0}$ and cavity parameters $G_{c}=0.2E_{d}$, $\kappa=5E_{d}$.
The results are shown in Fig. 3 (a), (b), where we compare the occupation of the excited electronic state $p_{e}$ and the cavity in the steady state for ten molecules and different coherent driving strengths $E_{d}$ (orange circles). The results are contrasted with a second set of BBGKY-HEOM simulations (black crosses), where the coherent drive and the vibrational bath have been combined into an effective incoherent pumping, such that the molecules evolve according to
$\dot{\rho}_{incoh}=$ $-i[H_{incoh}^{I},\rho_{incoh}]+\Gamma_{\downarrow}\sum_{j}\mathcal{D}[\sigma_{-}^{j}](\rho_{incoh})$
$+E_{d}\sum_{j}\mathcal{D}[\sigma_{+}^{j}](\rho_{incoh}),$
$H_{incoh}^{I}=$ $\sum_{\lambda}g_{\lambda}^{cav}\left(\sum_{j}\sigma_{+}^{j}a_{\lambda}+\sum_{j}\sigma_{-}^{j}a_{\lambda}^{\dagger}\right)$
$+\sum_{\lambda}(\omega_{\lambda}^{cav}-\omega_{0})a_{\lambda}^{\dagger}a_{\lambda}.$ (21)
We focus on the qualitative behavior as coherent driving and incoherent pumping will result in different steady-states for a given value. Fig. 3 (a) demonstrates the required population inversion for a lasing transition. In contrast to the predictions from the incoherent drive, the full model does not indefinitely increase its inversion, but rather reaches a plateau at large driving strengths. There the coherent drive dominates the dynamics, leading to 0.5 inversion. Interestingly, we find a maximal inversion at $E_{d}\approx 0.3E_{max}$, with the corresponding time evolution of $p_{e}$ shown in the inset. For the steady state values of the cavity occupation in Fig. 3 (b) we find that the results for the full model feature distinct maxima on top of a background that looks similar to the results obtained from the incoherent drive. These maxima become even more pronounced for larger numbers of molecules $N$, as they are enhanced by the self-amplifying lasing process, see Fig. 3 (c).
Let us rationalize this effect by starting from a single dissipation-less vibrational mode for the molecule
$H_{simple}=E_{d}\sigma_{x}+g^{vib}\sigma_{+}\sigma_{-}(b+b^{\dagger})+\omega^{vib}b^{\dagger}b.$ (22)
Looking at Fig. 2 (c) we can see that the four level picture breaks down if the two equilibrium positions are so close that $\lvert\innerproduct{4}{3}\rvert\approx 1$. In that case the molecule would look more like a two-level system in which a population inversion is not possible by means of coherent driving. As $b+b^{\dagger}$ corresponds to the displacement from the ground state equilibrium position as defined in Eq. (III.2), a significant occupation of the vibrational modes is beneficial for the lasing process. In Appendix G we show that a resonance occurs if the driving strength $E_{d}$ is equal to $\omega^{vib}/2$, resulting in a maximized occupation of the vibrational mode. In fact, we find roughly $\langle b\rangle\propto 1/(\omega^{vib}-2E_{d})$. Thus, the displacement of the vibrational mode is enhanced at certain frequencies, leading to a more efficient lasing process as explained above. Indeed, comparing Fig. 3 (c) to Fig. 2 (d) we find that the maxima of $\langle a^{\dagger}a\rangle$ line up with the peaks in the spectral density according to $\omega^{peak}=2E_{d}^{peak}$. Furthermore, artificially removing the peak at $\omega\approx 1.3E_{max}$ in the spectral density (Fig. 2 (d)) we find that also the 2nd maxima in the cavity occupation vanishes, as shown in Appendix G.
The many-body picture employed in this section provides straight-forward access to vibronic coupling and recommends itself as a natural basis to describe few-emitter lasing. It should be noted that plasmonic surfaces are prone to strong nonradiative losses via Landau damping and subsequent hot-carrier generation, beneficially utilized in plasmonic catalysis [84, 85, 86]. Future work should therefore consider exploring the impact of electronic and vibrational heating on the optical characteristics of few-emitter lasing [87].